• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Mina Kat

    Log entry time 00:04:04 on April15,2007

    Entry number 199080

    keyword=Shift Summary

    Continuing Q**2= 55 fm-2 kinematic point.

    Frequent beam trips.
    18:10 Called MCC: problem not current related, but due to a cavity
    MCC is working on problem

    18:54 MCC rew Chief called to inform us that the attenuator is rising, It will be maxed out around midnight. At approx. that time
    the max current we can get will be ~80uA. MCC will do a spot move tomorrow morning. That will get us to the current value we are getting now (~95uA).

    19:46 VESDA Beam Way alert, reset and now OK

    20:20 Left Q1 trip, Powered on, polarity negative, OK now

    Frequent beam trips again. Ended run 2476

    21:51 MCC Crew Chief called to tell us they will be checking steering in the injector to fix problem. No beam

    22:00 Beam back at ~91uA

    23:27 Left Q1 trip, ended run, powered on, OK now

    Run Summary:

    2466-2472, 1.4M at a current of 95uA
    2473 0.35M Left Q1 trip, ended run
    2472-2475, 1.4M at a current of 95uA
    2476 0.6M, ended run due to frequent beam trips
    2477-2478, 1.4M at a current of 90uA
    2479, 0.84 Left Q1 trip, ended run
    2480 in progress